PLEASE HELP!!!
It is known that a certain prescription drug produces undesirable side effects in 30% of all patients who use it. Among a random sample of 8 patients using the drug, find the probability of each event. Round each answer to the nearest hundredth.
a) None have undesirable side effects
b) Exactly one has undesirable side effects
c) Exactly two have undesirable side effects
d) More than two have undesirable side effects

Answer by stanbon(75887)

It is known that a certain prescription drug produces undesirable side effects in 30% of all patients who use it. Among a random sample of 8 patients using the drug, find the probability of each event. Round each answer to the nearest hundredth.
------
Binomial Problems with n = 8 ; p(side effects) = 0.3 ; p(no side) = 0.7
--------------------------------------------------
a) None have undesirable side effects
P(8 have no side) = 0.7^8 = 0.0576
-----------------------------------
b) Exactly one has undesirable side effects
P(x = 1) = 8C1*0.3*0.7^7 = 8*0.3*0.0824 = 0.1977
---------------------------------------------------
c) Exactly two have undesirable side effects
P(x = 2) = 8C2*0.3^2*0.7^6 = 28*0.09*0.1176 = 0.2965
---------------------------------------------------
d) More than two have undesirable side effects
p(3<= x <=8) = 1 - binomcdf(8,0.3,2) = 0.4482
=========================
Cheers,
Stan H.
=========================